# On the Privacy of Selection Mechanisms with Gaussian Noise

### v1

@@ 2 PRELIMINARIES
$\hat{\rho}=\rho+\mathcal{N}(0,\sigma_{X}^{2})$
for $t=1,2,\ldots$ do
 $\hat{q_{t}}=q_{t}(D)+\mathcal{N}(0,\sigma_{Z}^{2})$
 if $\hat{q_{t}}\geq\hat{\rho}$ then
@@ 3.1 Warm-Up: Gaussian Report Noisy Max
Let $M$ be the Gaussian Report Noisy Max mechanism with standard deviation $\sigma$ applied to $d>1$ bounded queries $q_{1},\ldots,q_{d}:\mathcal{X}\to[a,b]$, each with sensitivity bounded by $\Delta$. Let $c=b-a$. Let $\Phi(\cdot)$ be the standard Gaussian CDF. Then $M$ satisfies $\epsilon$-DP with
@@ 3.2 Ex-Post Gaussian Above Threshold
Consider a function ${}\epsilon_{\text{p}}:\cal{O}\to\mathbb{R}_{+}\cup\{\infty\}$. A randomized mechanism $M:\mathcal{X}\to\cal{O}$ satisfies $\epsilon_{\text{p}}$-ex-post-DP if for any possible output $o\in\cal{O}$ and any pair of neighboring datasets we have $\mathcal{L}_{M,D,D^{\prime}}(o)\leq\epsilon_{\text{p}}(o)$.
@@
Given a stream $q_{1},q_{2},\ldots:\mathcal{X}\to[a,b]$ with global sensitivity $\Delta$, the Gaussian Above Threshold mechanism (Algorithm 3), halting at time step $t$ with $o=\{\bot^{t-1}\top\}$, satisfies $\epsilon_{\text{p}}$-ex-post-DP with
$\epsilon_{\text{p}}(o)$ $=\frac{\mathbb{E}_{x\sim\mathcal{N}(0,1)}\left[\psi_{\Delta}(\sigma_{X}x)^{(t-1)}\cdot\beta_{\Delta}(\sigma_{X}x)\right]}{\mathbb{E}_{x\sim\mathcal{N}(0,1)}\left[\psi_{0}(\sigma_{X}x)^{(t-1)}\cdot\beta_{0}(\sigma_{X}x)\right]}\enspace.$
@@ 3.3 Fully-Adaptive Composition with Ex-Post Privacy Guarantees
    $o_{t}=M_{t}(D;o_{1:t-1})$
    $\epsilon_{t}=\epsilon_{\text{p},t}(o_{t})$
     Release $o_{t}$
@@
Suppose the mechanisms $M_{1},M_{2},\ldots$ provided to Algorithm 4 are such that $M_{t}$ is $(\epsilon_{\max,t},\delta)$-pDP and $\epsilon_{\text{p},t}$-ex-post-DP for all $t$. Then Algorithm 4 satisfies $(\epsilon,\delta)$-DP.
@@ 3.4 Numerical Computation of Bounds
Figure 4: Scatter plot indicating the accuracy for UCI Bikes with $\rho=0.575$. and final privacy spend over a range of noise multipliers. Final privacy loss ($\epsilon$) is reported for FSRC (green). Threshold noise, $\sigma_{X}$ in the range of [0.09,0.15][0.09,0.15][ 0.09 , 0.15 ]. A clear separation in privacy accounting occurs over a range of noise multipliers.
@@ 4 EMPIRICAL RESULTS
Figure 5: Scatter plot indicating the accuracy and final privacy spend over a range of noise multipliers for $\rho=0.35$ with the LCL London Energy dataset. Privacy loss ($\epsilon$) is reported for FSRC (green). Threshold noise, $\sigma_{X}$, was evaluated in the range of [0.06,0.15][0.06,0.15][ 0.06 , 0.15 ]. Our accounting method provides benefits when $\sigma_{X}=0.06$.
@@ A.1 Pure DP Gaussian Report Noisy Max
[…]
We say that a random variable $Z$ has tails majorized by an exponential decay if the cumulative distribution function $F_{Z}$ of $Z$ is such that there exist positive constants $c_{1},c_{2}$ satisfying $F_{Z}(z)=O(e^{c_{1}z})$ for $z\to-\infty$ and $1-F_{Z}(z)=O(e^{-c_{2}z})$ for $z\to\infty$. This is a well-known sufficient condition for the cumulant generating function $K_{Z}(s)=\log\mathbb{E}_{\,}\left[e^{sZ}\right]$ to exist.
@@
Let $\xi(s)=\mathbb{E}_{\,}\left[e^{sZ}\right]$ be the moment generating function of $Z$. Taking the derivative of $R$ we get:
@@ A.2 Pure DP Above Threshold
$\hat{\rho}=\rho+\mathcal{N}(0,\sigma_{X}^{2})$
for $t=1,2,\ldots$ do
 $\hat{q_{t}}=q_{t}(D)+\mathcal{N}(0,\sigma_{Z}^{2})$
 if $\hat{q_{t}}\geq\hat{\rho}$ then
@@
Given a stream $q_{1},q_{2},\ldots:\mathcal{X}\to[a,b]$ with global sensitivity $\Delta$, the Gaussian Above Threshold mechanism (Algorithm 3) satisfies $\epsilon_{\text{post}}$-ex-post-DP with
$\epsilon_{\text{post}}(\{\bot^{t-1}\top\})$ $=\frac{\mathbb{E}_{x\sim\mathcal{N}(0,1)}\left[\Phi\left(\frac{\sigma_{X}x+\rho-(b-a)+\Delta}{\sigma_{Z}}\right)^{(t-1)}\Phi\left(\frac{-\sigma_{X}x-\rho+a+\Delta}{\sigma_{Z}}\right)\right]}{\mathbb{E}_{x\sim\mathcal{N}(0,1)}\left[\Phi\left(\frac{\sigma_{X}x+\rho-(b-a)}{\sigma_{Z}}\right)^{(t-1)}\Phi\left(\frac{-\sigma_{X}x-\rho+a}{\sigma_{Z}}\right)\right]}\enspace.$ (25)
@@
[…]
@@ A.3 Filtered Self-Reporting Composition
Suppose the mechanisms $M_{1},M_{2},\ldots$ provided to Algorithm 4 are such $M_{t}$ is $(\epsilon_{\max,t},\delta)$-pDP and $\epsilon_{\text{post},t}$-ex-post-DP for all $t$. Then Algorithm 4 satisfies $(\epsilon,\delta)$-DP.
@@
    $o_{t}=M_{t}(D;o_{1:t-1})$
    $\epsilon_{t}=\epsilon_{\text{post},t}(o_{t})$
     Release $o_{t}$
@@ B.3 Offline Selection with Gaussian Report Noisy Max
[…]
@@
(b) Comparison for $d=829$ on the LCL Energy Dataset.
@@ Appendix C COMPARISON WITH PRIVACY FILTERS
 $o_{t}=M_{t}(D)$
 if ${\cal{F}}_{\epsilon_{\text{max}},\delta_{\text{max}}}((\epsilon_{t})_{t\geq 1},(\delta)_{t\geq 1})=\text{HALT}$ then
@@ Appendix D PURE DP ONLINE SELECTION: ADDITIONAL EXPERIMENTS
Figure 17: Scatter plot indicating the accuracy and final privacy spend over a range of noise multipliers for thresholds with the UCI Bikes dataset. Final privacy loss ($\epsilon$) is reported for Filtered Self-Reporting Composition (green). Threshold noise, $\sigma_{X}$, evaluated in the range of [0.09,0.15][0.09,0.15][ 0.09 , 0.15 ]. We note a clear separation in privacy accounting over a range of noise multipliers.
@@
London Energy Dataset
We plot Gaussian Above Threshold under the same utilty metric on the LCL London energy dataset, with 5,564 customers. In Fig. 19 we plot a range of calibrations. As the threshold decreases, we witness more queries released and therefore a greater privacy spend. we show the effect over time in Fig. 20.
Figure 19: Scatter plot indicating the accuracy and final privacy spend over a range of noise multipliers for thresholds with the LCL London dataset. Final privacy loss ($\epsilon$) is reported for FSRC (green). Threshold noise, $\sigma_{X}$, was evaluated in the range of [0.06,0.15][0.06,0.15][ 0.06 , 0.15 ]. With this dataset, our accounting method provides benefits when $\sigma_{X}>0.06$.

### v2

@@ 2 PRELIMINARIES
$\hat{\rho}=\rho+\mathcal{N}(0,\sigma_{X}^{2})$ for $t=1,2,\ldots$ do
       $\hat{q_{t}}=q_{t}(D)+\mathcal{N}(0,\sigma_{Z}^{2})$ if $\hat{q_{t}}\geq\hat{\rho}$ then
@@ 3.1 Warm-Up: Gaussian Report Noisy Max
Let $M$ be the Gaussian Report Noisy Max mechanism with standard deviation $\sigma$ applied to $d>1$ bounded queries $q_{1},\ldots,q_{d}:\mathcal{X}\to[a,b]$, each with sensitivity bounded by $\Delta$. Let $c=b-a$. Let Φ()normal-$\Phi(\cdot)$ be the standard Gaussian CDF. Then $M$ satisfies $\epsilon$-DP with
@@ 3.2 Ex-Post Gaussian Above Threshold
[…]
@@
[…]
@@ 3.3 Fully-Adaptive Composition with Ex-Post Privacy Guarantees
             $o_{t}=M_{t}(D;o_{1:t-1})$ $\epsilon_{t}=\epsilon_{\text{p},t}(o_{t})$ Release $o_{t}$
@@
[…]
@@ 3.4 Numerical Computation of Bounds
Figure 4: Scatter plot indicating the accuracy for UCI Bikes with $\rho=0.575$. and final privacy spend over a range of noise multipliers. Final privacy loss ($\epsilon$) is reported for FSRC (green). Threshold noise, $\sigma_{X}$ in the range of $[0.09,0.15]$. A clear separation in privacy accounting occurs over a range of noise multipliers.
@@ 4 EMPIRICAL RESULTS
Figure 5: Scatter plot indicating the accuracy and final privacy spend over a range of noise multipliers for $\rho=0.33$ with the LCL London Energy dataset. Privacy loss ($\epsilon$) is reported for FSRC (green). Threshold noise, $\sigma_{X}$, was evaluated in the range of $[0.04,0.16]$. Our accounting method provides benefits when $\sigma_{X}=0.04$.
@@ A.1 Pure DP Gaussian Report Noisy Max
Let $-\infty\leq a<b\leq\infty$ and let $f,g:[a,b]\to\mathbb{R}$ be continuous differentiable functions on $(a,b)$, with $f(a)=g(a)=0$ or $f(b)=g(b)=0$, and $g^{\prime}(x)\neq 0$ for $x\in(a,b)$. If $f^{\prime}/g^{\prime}$ is non-decreasing in $(a,b)$, then so is $f/g$.
We say that a random variable $Z$ has tails majorized by an exponential decay if the cumulative distribution function $F_{Z}$ of $Z$ is such that there exist positive constants $c_{1},c_{2}$ satisfying $F_{Z}(z)=O(e^{c_{1}z})$ for $z\to-\infty$ and $1-F_{Z}(z)=O(e^{-c_{2}z})$ for $z\to\infty$. This is a well-known sufficient condition for the cumulant generating function $K_{Z}(s)=\log\mathbb{E}\left[e^{sZ}\right]$ to exist.
@@
Let $\xi(s)=\mathbb{E}\left[e^{sZ}\right]$ be the moment generating function of $Z$. Taking the derivative of $R$ we get:
@@ A.2 Pure DP Above Threshold
$\hat{\rho}=\rho+\mathcal{N}(0,\sigma_{X}^{2})$ for $t=1,2,\ldots$ do
       $\hat{q_{t}}=q_{t}(D)+\mathcal{N}(0,\sigma_{Z}^{2})$ if $\hat{q_{t}}\geq\hat{\rho}$ then
@@
[…]
@@
Now note that in the bound we set $q_{i}-q_{i}^{\prime}=-\Delta$ for $i<t$ and $q_{t}-q_{t}^{\prime}=\Delta$. Since all the queries are bounded in $[a,b]$ we have that $y_{i}=q_{i}^{\prime}=q_{i}+\Delta\in[a+\Delta,b]$ for $i<t$ and $y_{t}=q_{t}^{\prime}=q_{t}-\Delta\in[a,b-\Delta]$. Taking the supremum over these ranges completes the proof. ∎
@@ A.3 Filtered Self-Reporting Composition
[…]
@@
             $o_{t}=M_{t}(D;o_{1:t-1})$ $\epsilon_{t}=\epsilon_{\text{post},t}(o_{t})$ Release $o_{t}$
@@ B.3 Offline Selection with Gaussian Report Noisy Max
Figure 9: Privacy / utility when performing Gaussian Report Noisy Max on a query over $d=364$. $90\%$ Accuracy is attainable with our method whereas the same would only be possible with over double the privacy budget.
@@
(b) Comparison for $d=365$ on the LCL Energy Dataset.
@@ Appendix C COMPARISON WITH PRIVACY FILTERS
[…]
@@ Appendix D PURE DP ONLINE SELECTION: ADDITIONAL EXPERIMENTS
Figure 17: Scatter plot indicating the accuracy and final privacy spend over a range of noise multipliers for thresholds with the UCI Bikes dataset. Final privacy loss ($\epsilon$) is reported for Filtered Self-Reporting Composition (green). Threshold noise, $\sigma_{X}$, evaluated in the range of $[0.09,0.15]$. We note a clear separation in privacy accounting over a range of noise multipliers.
@@
London Energy Dataset.
We plot Gaussian Above Threshold under the same utility metric on the LCL London energy dataset, with 5,564 customers. In Fig. 19 we plot a range of calibrations. As the threshold decreases, we witness more queries released and therefore a greater privacy spend. we show the effect over time in Fig. 20.
Figure 19: Scatter plot indicating the accuracy and final privacy spend over a range of noise multipliers for thresholds with the LCL London dataset. Final privacy loss ($\epsilon$) is reported for FSRC (green). Threshold noise, $\sigma_{X}$, was evaluated in the range of $[0.04,0.16]$. With this dataset, our accounting method provides benefits when $\sigma_{X}>0.06$.